OpenStudy (ilovecomputers1):

Find the middle term of ( 2/x+x/2)^10

OpenStudy (anonymous):

\[\binom{10}{5}\left(\frac{2}{x}\right)^5\left(\frac{x}{2}\right)^5\] is the first step